OpenStudy (anonymous):

find the distance between (-7,5) and (-6,-4)

OpenStudy (anonymous):

that answer is not on my choices

OpenStudy (anonymous):

woooooh....distance formula is not that!

OpenStudy (anonymous):

distance formula is: SQRT((y2-y1)^2 + (x2-x1)^2))

OpenStudy (anonymous):

(-4-5)^2-(-6+4)^2=81-4=77----77\[\sqrt{77}\]

OpenStudy (anonymous):

NO. That is wrong.

OpenStudy (anonymous):

\[d = \sqrt{(x2-x1)^{2} + (y2-y1)^{2})}\]

OpenStudy (anonymous):

opss sorry (-6+7)^2=1 81-1=80 \[\sqrt{80}\]

OpenStudy (anonymous):